OpenStudy (anonymous):

1) If h(x)=x^2-5 find h(-3) 2) Given the function g(x)=x2+1, evaluate g(5x)) 3) Joe drives 75 miles per hour. Pete drives 70 miles per hour. Write a function that defines how much further Joe will drive after t hours. 4) If f(x)=2x-15, find f(-5) 5) Given the function h(r)=2r-5, evaluate h(r+5) 6) If f(x)=3x and g(x)=2-6, find f(x)*g(x) 7) If f(x)=5x-4 and g(x)=2x+2, find g(x)-f(x) 8) If p(r)=10-15 and q(r)=5, find p(r)/q(r) 9) Find the area of the triangle with base equal to (2x-3) and height equal to (4x+5). Area of the triangle equals 1/2b*h.

OpenStudy (anonymous):

in first put -3 in place of x and find the value

OpenStudy (anonymous):

in 2nd do the same,put 5x in equation

OpenStudy (anonymous):

3rd is arithmetic progression in 1 hour he will be 5 miles further so in t hours he will be 5t miles further

OpenStudy (anonymous):

4th you have put value of -5 and solve

OpenStudy (anonymous):

in 5th put r+5 instead of r and solve

OpenStudy (anonymous):

whats 6th? can you rewrite g(x)

OpenStudy (anonymous):

6) If f(x)=3x and g(x)=2-6, find f(x)*g(x)

OpenStudy (anonymous):

okay then you have to do it like this 3x(2-6) = 6x-18x=-12x

OpenStudy (anonymous):

7th is simple subtraction 5x-4-(2x+2) = 3x-6

OpenStudy (anonymous):

now p(r)/q(r) same put the values 10-15/5 = -5/5 = -1

OpenStudy (anonymous):

now they have given you formula in 9th one so area = 1/2 (2x-3)(4x+5) =1/2( 8x^2+10x-12x-15) = 1/2( 8x^2-2x-15) keep it like this as final answer

OpenStudy (anonymous):

Thanks so much! If I have any more questions I'll shoot you another message, thanks again!

OpenStudy (anonymous):

welcome :)